James is 5 times as old as Mike. In 9 years James will only be twice as old as Mike is then. How old is eache now?
My answer keeps coming out wrong, help please.!?

Answer by lwsshak3(11628)

James is 5 times as old as Mike. In 9 years James will only be twice as old as Mike is then. How old is eache now?
***
let x=mike's age
5x=james' age
..
In 9 yrs:
x+9=mike's age
5x+9=james' age
james is twice as old as mike
5x+9=2(x+9)
5x+9=2x+18
3x=9
x=3
5x=15
..
now:
mike's age=3
james' age=15